OpenStudy (lin.ivory):

log(3x+2)+log(x-1)=2 how to solve for x?

OpenStudy (anonymous):

(x-1 should be >0 and 3x+2 > 0 for logarithm to be defined) log((3x+2)(x-1))=2 log((3x^2-3x+2x-2))=2 taking out log 3x^2-x-2=10^2 3x^2-x-2=100 3x^2-x-102=0 facotring these shall give you (x-6)(3x+17)=0 x=6 or x=-17/3 From the first sentence, we can eliminate -17/3 making x=6 is the right answer

OpenStudy (lin.ivory):

Thank you!